

********************************************************************
UARS WEEKLY STATUS REPORT - 19 December 1997
********************************************************************
***********
SOLSTICE
***********
Happy Winter SOLSTICE and Merry Christmas!
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
Happy Holidays from the SUSIM team!
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE was operating successfully when turned off on December 15th. All
engineering parameters were nominal and the science data products continue
to be of excellent quality. HALOE is scheduled to be turned back on
December 29th.
***********
MLS
***********
The MLS instrument has been powered off since Dec 2. MLS is scheduled to
begin south viewing operations on Dec 22.
Testing has now been completed of the software running on the CDHF for
producing the upper troposphere humidity (UTH) product from MLS. Production of
catalogued data for the entire mission will now begin, and is expected to be
completed within a few months.
These UTH data are produced using existing MLS Version 4 radiances and tangent
pressure. They will be catalogued on the CDHF as "Version 4" in order to
maintain numbering consistency with the version of other MLS data from which
they are produced.
The retrieved quantities are relative humidity (with respect to ice) at UARS
pressure surfaces of 147, 215, 316 and 464 hPa. Level 3AT and Level 2 files
are produced. Retrieved relative humidity values greater than 100% indicate
the presence of cloud. The Level 2 UTH file has additional information, such
as coincident temperatures (from NCEP), and the corresponding water vapor
mixing ratios.
Our initial analyses indicate these data are accurate to around 20%, with the
464 hPa values being less reliable than those at higher altitudes. It should
be emphazed, however, that these data have not yet been validated and
validation activies will take place over the next several months, culminating
in a validation publication.
Any questions on these data should be referred to Dr. William G. Read
(telephone 818-354-6773; e-mail bill@mls.jpl.nasa.gov; fax 818-393-5065).
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF will be unavailable on Sunday 12/21/97, for an estimated two hour
window within the time period of 12:00 - 18:00 EST. A GSFC power outage in
building 1 will disable all network connectivity at GSFC during this time.
The CDHF will remain up, but unavailable via the network.
The CDHF has received a total of 353 CD-ROM requests as of 19-December-97.
There were no orders received. One prior order was shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess
25 days PEM (3.32) Prod (1350R-685C=665L)
The CDHF notified the DAAC that the October 1997 current day data has been
loaded to the DAAC transfer tables for HALOE, HRDI, PEM, NMC, and UKMO.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 December 1997 (Orbit 34072, GMT Day 339) through 12
December 1997 (Orbit 34177, GMT Day 346).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 340/0035 34073 On
WINDII 341/1145 34095 Off
PEM 341/2333 34102 ZEPS off
HRDI 342/0044 34103 On
HRDI 346/2319 34177 Off, low power
The beta angle is now decreasing toward a minimum angle of 0.0 degrees on 17
December 1997 (DOY 351).
Four additional SSPP Solar Offset experiments were performed this week - UTC
Day 340, 341, 342 and 345. These experiments were identical to the normal
solar offset that is performed on a weekly basis.
Changes to Telemetry Monitors (TMONs) 1 and 2 were uplinked on 11 December 1997
(Orbit 34158, 345/16:34:35 UTC) (see Other Significant Events).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.6 end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 degrees C. Additional battery performance data for each battery on
06 December 1997 and 12 December 1997 are provided in Appendix D.
The clock error ranged from -11.6 to +10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 06
% HALOE 01
% HRDI 02
% PEM 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 02
TRANSPONDER FREQUENCIES:
Transponder A: 97/343 - 2287.497007
Transponder B: 97/344 - 2287.499121
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 97/345 show the observatory data loss to be 20 hours, 41
minutes, 12.916 seconds (no change). This is a 0.0379 percent data loss which
equals a 99.9621 percent data capture for the mission.
Modifications to the OBCSTF resident POCC database were completed. The
database (Version 65) has been tested on the UTTS and will be officially
released next week. This Database includes changes in the battery voltage red
and yellow limits to reflect the two-battery operation of the spacecraft. The
new yellow and red limits are 23.68 V and 23.36 V respectively, which
correspond to the new TMON 1 and 2 limits.
The desired receive frequencies for the UARS transponders was transmitted by
memorandum to T. Marsh on 12 December 1997. The new receive frequencies for
transponder A and B will be 2106.40350 Mhz and 2106.40535 Mhz, respectively.
The FOT is still working out the details of participating in the IP return-link
telemetry test that the network is conducting.
Flight Dynamics Facility (FDF) continues the generation process for a new Star
Catalog to fix the invalid precession rate for star 48000001.
No testing of the new CMS software release (R7B1V2) occurred this week.
UTTS testing of the 10.08 POCC AP ground system software was conducted on 10
December 1997. Previous tests of the 10.08 ground system software revealed a
problem with the software that considerably extended the dump compare process
time to unacceptable levels. Test problems were encountered because OBC dumps
were not completing. Because of this, a dump file to use for the compare with
the ground image file could not be generated. The problem was believed to be
with the Q-channel line of the simulator configuration. However, the dump
compare portion of the process was conducted with the updated ground image file
and the last dump file on AP Pack PE 41. The dump compare time was
approximately one minute with an expected large amount of miscompares.
The test was continued on 12 December 1996 and a current dump file was directly
downlinked to pack PE 41 and compared with the associated reference file using
Version 10.08 of the POCC software. The command verification during this test
was also approximately one minute with no miscompares. POCC AP Software
Version 10.08 has corrected the dump compare excessive time-to-complete problem.
UARS Test and Training Simulator (UTTS) testing to correct the Q-channel line
problem have been scheduled and will be conducted this week.
New limits for ACRIM's ACHTRI2 and ACHTRV1 telemetry points were received from
the Principal Investigator. For ACHTRV1, the new red-low limit will be set at
0.5 V and the new yellow-low limit will be set at 0.6 V. For ACHTRI2, the new
red-low limit will be set at 1.15 A and the new yellow-low limit will be set at
1.2 A. These updated limits were received after the latest database was updated
and will be added to the LIMDEF procedure until a new database is released.
Operational printing of all CDHF-generated FOT telemetry analysis products from
the UMPGAL computer (new MPG Alpha machine) was implemented beginning on 10
December 1997. Shadow processing of the level-0 files was also begun on UMPGAL
as of 11 December 1997. There appears to be a memory allocation problem on
UMPGAL which prevents full completion of the level-0 processing on this
machine. SSAI is investigating the problem, and further tuning of the
computer operating system is expected to resolve this problem within the next
week.
OTHER SIGNIFICANT EVENTS
Successfully tested changes to TMONs 1 and 2 (with the UTTS) on 10 December
1997. Lowered the voltage trip limit for TMON-01 to 23.68 V (i.e., load bus
voltage, battery-2 voltage, and battery-3 voltage must be below 23.68 V for the
first step of load shedding to occur). Lowered the voltage trip limit for TMON-
02 to 23.36 V (i.e., load bus voltage must be below 23.36 V for the second step
of load shedding to occur.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 17 December 1997
at 311/132708.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
will probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
Frank Randall, Flight Operations Director (FOD), submitted his resignation
effective 08 December 1997 to accept a position at the John Hopkins Applied
Physics Laboratory (APL).
Dave Stovall, Evaluation/Support Engineer, has been assigned as a Flight
Operations Director.
Dimitrios Mantziaras, recent Aerospace Engineer graduate of Virginia Tech,
has accepted a position as Evaluation/Support Engineer.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 December 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34073 340/0035 97-611 HALOE Power ON
34076 340/0533 97-612 OBC/CLK Clk Adj fr 0.00 to 20.6 msec/day
34102 341/2333 97-614 PEM Power Off ZEPS
34095 341/1145 97-613 WINDII Power OFF
34103 342/0044 97-615 HRDI Power ON
34118 343/0045 97-616 OBC Clk Adj fr 20.6 to 0.00 msec/day
34129 343/1826 97-617 OBC Disabled TMONs 45 & 47 (statbuf)
34151 345/0556 97-618 OBC Clk Adj fr 0.00 to 20.6 msec/day
34158 345/1626 97-619 OBC Changed TMON 1 & 2 Voltage Limits
34173 346/1700 97-620 OBC Clk Adj fr 20.6 to 0.00 msec/day
34177 346/2319 97-621 HRDI "Power OFF, Low pwr mode"
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 December 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSSOC2 Y-LO Battery 2 Current SOC 34079,86,94,99-101,108-09,
124-25,133-46,153-62,169
OBSVLBEON Y-LO Load Bus Volts at EON 34147,63,65-69
OBSBATVEON2 Y-LO Battery 2 voltage at EON 34147,165-77
OBSBATVEON3 Y-LO Battery 3 voltage at EON 34147,63,65-77
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WISRCLSR WINDII Laser is Operating 3409
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 December 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34076 340/05:37 (06 Dec 97) from 00.0 to 20.6 msec/day
34118 343/00:50 (09 Dec 97) from 20.6 to 00.0 msec/day
34151 345/06:04 (11 Dec 97) from 00.0 to 20.6 msec/day
34173 346/17:06 (12 Dec 97) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 December 1997
06 Dec 1997 - GMT Day 340 Beta = 47.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.36/-4.94 +0.0/ +0.0 18.7 0.000 0.0
2 +33.6/ -16.8 +5.47/+3.87 +28.8/-16.0 25.9 1.024 21.5
3 +84.0/ -11.2 +4.51/+2.60 +27.6/-15.6 25.8 1.024 20.8
12 Dec 1997 - GMT Day 346 Beta = 22.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.7 0.000 0.0
2 +22.4/ -16.8 +6.44/+5.15 +32.0/-15.2 24.8 1.021 23.1
3 +112.0/ +0.0 +5.15/+3.87 +31.2/-16.0 24.8 1.022 22.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 December 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-126 FOT 34104 342/023130 N/A TAC Crashed
97-124 FOT 34123 343/0809 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-125 FOT 34029 337/02:39 KCRTUS NON-OP
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
SEVNTFW50